What is the mole fraction of sucrose (C12H22O11, molar mass = 342 g/ mol) and water ( molar mass = 16 g / mol ) in 5.25% by mass sucrose?

A.0.0035, 0.9965

B.0.0021, 0.9979

C.0.0029, 0.9971

D.0.0018, 0.9982



1
Expert's answer
2021-10-24T01:28:20-0400

513 grams of sucrose have moles = 513 / 342 = 1.5 moles .

t2liters of water have 2000 ml of water

density of water is 1 g /

so , mass of water = 2000

moles of water = 2000 / 18 = 111.11 m

so , mole fraction of sucrose = moles of sucrose / total moes of solution


= 1.5 / 1.5 + 111.= 0.0133

mole fraction of water = 1 - mole fraction of sucrose = 1- 0.013320 = 0.9866720 .11on .ol . g .ml .te . sucrose = 1- 0.013320 = 0.98667